# Fabian Mauthe

### v1

@@
Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Hagmeyer, S.; Mauthe, F.; Zeiler, P. (2021): Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087.
@@ 1 Prognostics and Health Management
In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview of publicly available degradation data sets is presented as well as briefly described. The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
@@ 2 Tasks of Diagnostics and Prognostics
The prognosis task poses the greatest challenge in diagnostics and prognostics within PHM [6]. This difficulty stems from the intricate nature of failure mechanisms and the multitude of factors influencing degradation progression. Therefore, sufficient degradation data is of huge importance for this task and needs to be comprehensively available.
@@ 3 Data Sets for Diagnostics and Prognostics
Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically. This is the extended overview from [4], which already contains 70 data sets.
@@
  • Domain and Application: The associated domain from which the data set originates and the respective application within this domain.
@@
4 Description of Data Sets
The overview of publicly available degradation data sets in Table Appendix: Overview of Publicly Available Data Sets comprises a total of 79 data sets. These data sets originate from eleven different domains. The domains as well as the applications presented do not claim to be exhaustive and are based solely on the available data sets. With 17 assigned data sets, the Mechanical component domain is the most extensive, followed by the Electrical component domain with 13 assigned data sets. The two most frequently occurring applications, Bearing (ten) and Battery (six), also originate from these two domains. In total, the domains contain 36 different applications.
Except for the Building domain, which is sparsely represented in PHM, the data sets are evenly distributed for the remaining domains, with at least four and at most nine assignments per domain. With only one data set for each application, the Mechatronic system domain contains the greatest variation of applications. Moreover, there are five data sets with unknown applications and therefore, no domain can be assigned. Data sets with degradation data resulting from a simulated application are assigned to the corresponding domains but marked. With 32 assigned data sets for prognosis and 29 for diagnosis, these tasks are the most represented. For fault detection, 15 data sets are assigned, and for health assessment, 3 data sets are assigned. The Electrical component domain contains the most data sets (ten) for the task prognosis, followed by the Drive technology domain with six and the Manufacturing process as well as Process technology domains with four assigned data sets each.
Of the 17 assigned data sets in the Mechanical component domain, only two are for task prognosis. However, for the task diagnosis, the Mechanical component domain contains the most data sets with eleven assignments. Followed by the Mechatronic system domain with five and the Robotic domain with four assigned data sets. Regarding fault detection, the domains Mechanical component and Production system contain the most data sets with four assignments each. With two data sets with an Unknown domain and one data set from the Drive technology domain, the overview contains the fewest data sets for the health assessment task.
@@ 4 Description of Data Sets
Overview of publicly available data sets
No.
Data Set Designation
Reference to Data Origin
Domain
Application
Task URL
@@
No.
Data Set Designation
Reference to Data Origin
Domain
Application
Task URL
@@
1
AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309.
Mechatronic system
@@
Diagnosis
Link
2
AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
Manufacturing process
@@
Diagnosis
Link
3
Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France).
Unknown
@@
Fault detection
Link
4
Backblaze - Hard Drive Stats
Backblaze, Inc.
Drive technology
@@
Fault detection
Link
5
Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA
Mechanical component
@@
Diagnosis
Link
6
GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191
Simulation
@@
Fault detection
Link
7
Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
Mechatronic system
@@
Fault detection
Link
8
Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
Manufacturing process
@@
Fault detection
Link
9
Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
@@
Prognosis
Link
10
Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom
Manufacturing process
@@
Prognosis
Link
11
Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Robotic
@@
Diagnosis
Link
12
Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance
Unknown
@@
Fault detection
Link
13
Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
@@
Fault detection
Link
14
Kaggle - Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi
Production system
@@
Fault detection
Link
15
Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
@@
Prognosis
Link
16
Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRP-CMS
Production system
@@
Fault detection
Link
17
Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
@@
Prognosis
Link
18
Kaggle - Pump
Kaggle User: UnknownClass
Process technology
@@
Fault detection
Link
19
Kaggle - Sensor Fault Detection
Schneider-Electric, France
Electrical component
@@
Fault detection
Link
20
Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
@@
Fault detection
Link
21
Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2
Drive technology
@@
Prognosis
Link
22
Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
Drive technology
@@
Diagnosis
Link
23
Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
Mechanical component
@@
Diagnosis
Link
24
MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
Mechanical component
@@
Diagnosis
Link
25
NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proccedings of the Annual Conference of the PHM Society, 15(1). https://doi.org/10.36001/phmconf.2023.v15i1.3463

### v2

@@
Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Fabian Mauthe, Luca Steinmann, and Peter Zeiler. Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management. In Eirik Bjorheim Abrahamsen, Terje Aven, Frederic Bouder, Roger Flage, and Marja Yloenen, editors, 35th European Safety and Reliability Conference. (Submitted). Research Publishing, 2025.
@@ 1 Prognostics and Health Management
In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview from [5] of publicly available degradation data sets is presented. The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
@@ 3 Data Sets for Diagnostics and Prognostics
Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically. This is the extended overview from [4], presented in [5].
@@
  • Domain, Application, and Signals: The associated domain from which the data set originates and the respective application within this domain. In addition, the signals present in the respective data set are listed.
@@
The detailed PHM-specific analysis of the data sets can be found in [5].
@@
{xltabular}
|p0.6cm|X|p4.1cm|p2.3cm|p1.0cm| Overview of publicly available data sets
No. Data Set Designation
Reference to Data Origin Domain
Application and Signal Task URL
@@
No. Data Set Designation
Reference to Data Origin Domain
Application and Signal Task URL
@@
1 4TU - Lifecycle ageing tests on commercial 18650 Li ion cell
Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
2 4TU - Motor Current and Vibration Monitoring Dataset
Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump. doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3 Process technology
Centrifugal pumps
Signals: current, vibration, voltage Diagnosis Link
3 AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309. Mechatronic system
@@ 3 AIDAR Lab - Air Compressor
Signals: acoustic emission Diagnosis Link
4 AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26. Manufacturing process
@@ 4 AIDAR Lab - Drill Bit
Signals: vibration Diagnosis Link
5 Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France). Unknown
@@ 5 Aramis - Data Challenge ESREL2020PSAM15
Signals: vibration Fault detection Link
6 Backblaze - Hard Drive Stats
Backblaze, Inc. Drive technology
@@ 6 Backblaze - Hard Drive Stats
Signals: temperature Fault detection Link
7 Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA Mechanical component
@@ 7 Bearing Data Center - Fan and Bearing
Signals: acoustic emission Diagnosis Link
8 Calce - Battery Data Repository
CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
9 GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191 Simulation
@@ 9 GitHub - Predictive Maintenance using PySpark
Signals: unknown Fault detection Link
10 GitHub - XJTU-SY Bearing Datasets
Biao Wang, Yaguo Lei, Naipeng Li, Ningbo Li (2020). A Hybrid Prognostics Approach for Estimating Remaining Useful Life of Rolling Element Bearings, IEEE Transactions on Reliability, vol. 69, no. 1, pp. 401-412. DOI: 10.1109/TR.2018.2882682. Mechanical component
Bearing
Signals: vibration Prognosis Link
11 Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm Mechatronic system
@@ 11 Kaggle - Air pressure system failures in Scania trucks
Signals: anonymized Fault detection Link
12 Kaggle - Bearings with Varying Degradation Behaviors
Mauthe, F.; Hagmeyer, S.; Zeiler, P. (2025). Holistic simulation model of the temporal degradation of rolling bearings. In E. B. Abrahamsen, T. Aven, F. Bouder, R. Flage, and M. Yloenen (Eds.), Proceedings of the 35nd European Safety and Reliability Conference (Submitted). Research Publishing. Mechanical component
Bearing (simulated)
Signals: vibration Prognosis Link
13 Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan Manufacturing process
@@ 13 Kaggle - CNC Mill Tool Wear
Signals: acceleration, current, feedrate, position, speed, voltage Fault detection Link
14 Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
@@ 14 Kaggle - Condition Data with Random Recording Time
Signals: flow rate, pressure Prognosis Link
15 Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom Manufacturing process
@@ 15 Kaggle - E-coating ultrafiltration maintenance
Signals: flow rate, pressure, temperature Prognosis Link
16 Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Robotic
@@ 16 Kaggle - Genesis Demonstrator
Signals: acceleration, current, force, speed Diagnosis Link
17 Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance Unknown
@@ 17 Kaggle - Microsoft Azure Predictive Maintenance
Signals: pressure, speed, vibration, voltage Prognosis Link
18 Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
@@ 18 Kaggle - One Year Industrial Component Degradation
Signals: position, speed, torque Fault detection Link
19 Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi Production system
@@ 19 Kaggle - Predictive Maintenance 1
Signals: unknown Fault detection Link
20 Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
@@ 20 Kaggle - Preventive to Predictive Maintenance
Signals: flow rate, pressure Prognosis Link
21 Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS Production system
@@ 21 Kaggle - Production Plant Data for Condition Monitoring
Signals: anonymized Fault detection Link
22 Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
@@ 22 Kaggle - Prognosis based on Varying Data Quality
Signals: dust feed, flow rate, pressure Prognosis Link
23 Kaggle - Pump
Kaggle User: UnknownClass Process technology
@@ 23 Kaggle - Pump
Signals: anonymized Fault detection Link
24 Kaggle - Sensor Fault Detection
Schneider-Electric, France Electrical component
@@ 24 Kaggle - Sensor Fault Detection
Signals: temperature, voltage Fault detection Link
25 Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
@@ 25 Kaggle - Versatile Production System
Signals: control parameters Fault detection Link
26 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
27 Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2 Drive technology
@@ 27 Mendeley - Brushless DC motor
Signals: current, position, temperature Prognosis Link
28 Mendeley - Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells
Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1 Electrical component
Battery
Signals: capacity Prognosis Link
29 Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil Drive technology
@@ 29 Mendeley - Diesel Engine Faults
Signals: pressure, temperature, vibration Diagnosis Link
30 Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3 Mechanical component
@@ 30 Mendeley - HUST Bearing
Signals: vibration Diagnosis Link
31 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1 Electrical component
Fuel Cell
Signals: unknown Prognosis Link
32 Mendeley - NMC cell 2600 mAh cyclic aging data
Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1 Electrical component
Battery
Signals: current, temperature Prognosis Link
33 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
Gabrielli, Alberto; Battarra, Mattia; Mucchi, Emiliano; Dalpiaz, Giorgio (2024). Physics-based prognostics of rolling-element bearings: The equivalent damaged volume algorithm, Mechanical Systems and Signal Processing, Volume 215, 2024, 111435, doi: 10.1016/j.ymssp.2024.111435. Mechanical component
@@ 33 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
Signals: vibration Prognosis Link
34 MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems Mechanical component
Bearing
Signals: vibration Diagnosis Link
35 NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proccedings of the Annual Conference of the PHM Society, 15(1). doi: 10.36001/phmconf.2023.v15i1.3463 Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
36 NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. ”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Mechanical component
Bearing
Signals: vibration Prognosis Link
37 NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Electrolytic Capacitors
Signals: impedance, voltage Prognosis Link
38 NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Electrolytic Capacitor
Signals: capacity, resistance Prognosis Link
39 NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Materials
Carbon Fibre-Reinforced Polymer Composite
Signals: operating condition, strain, vibration, x-ray Prognosis Link
40 NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
41 NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Transistor
Signals: , current, temperature, voltage Prognosis Link
42 NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
43 NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Manufacturing process
Milling
Signals: acoustic emission, current, vibration Prognosis Link
44 NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Transistor
Signals: current, temperature, voltage Prognosis Link
45 NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Prognosis Link
46 NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
Battery
Signals: current, temperature, voltage Diagnosis Link
47 NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
48 NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
Aircraft Engine*
Signals: anonymized, operating condition Prognosis Link
49 NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division Robotic
Industrial Robot
Signals: current, position, speed Diagnosis Link
50 OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019) Building
Buildings
Signals: pressure, speed, temperature Diagnosis Link
51 OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory Mechanical component
Gear
Signals: vibration Fault detection Link
52 Oxford - Oxford Battery Degradation Dataset
Birkl, C. (2017). Oxford Battery Degradation Dataset 1. University of Oxford. DOI: 10.5287/bodleian:KO2kdmYGg